OVHcloud Bare Metal Cloud Status

Current status
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#12571 — Vps Cloud - SBG
Incident Report for Bare Metal Cloud
We are having problems of communication with a switch. We are intervening.


Date: 2015-02-09 09:02:08 UTC
The new switch is up. VMs have all been switched over. Everything is configured.

Date: 2015-02-09 08:59:43 UTC
We've nearly completed the operation. We're currently verifying all the ACLs.

Date: 2015-02-09 08:58:41 UTC
We're still deploying a new switch. The operation is taking longer than anticipated.

Date: 2015-02-09 08:56:06 UTC
Currently the robots are disabled on the cluster.

Date: 2015-02-09 08:55:05 UTC
All the VMs are rebooting.
The deployment isn't fully complete. We will continue the reinstallation.

Date: 2015-02-09 08:52:50 UTC
There's no progress with the reconstruction. We will deploy a new switch.

Date: 2015-02-09 08:51:28 UTC
Currently resynchronising the master and secondary switch

Date: 2015-02-07 13:58:24 UTC
Some VPS has still network problems. We are searching for the cause.

Date: 2015-02-07 13:57:53 UTC
We restart the vcenter
Posted Feb 07, 2015 - 11:50 UTC